OpenStudy (anonymous):

the condition that the roots of px^2 - px + q = 0 are in the ratio p:q is?

OpenStudy (maheshmeghwal9):

Did sum of the roots & product of the roots help?

OpenStudy (maheshmeghwal9):

I mean did u try?

OpenStudy (anonymous):

no idea

mathslover (mathslover):

\[\huge{ax^2+bx+c=0}\] this is the general form of a quadratic equation in this quadratic equation \[\huge{x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}}\]

mathslover (mathslover):

now compare quadratic equation with given equation

mathslover (mathslover):

u will get that : a = p , b = - p and c =q right ?

OpenStudy (shubhamsrg):

let roots be a and b sum of roots = a+b = 1 and ab = q/p or 1/ab = p/q and you want a/b = p/q so 1/ab = a/b => a=1 or -1 (assuming one of the roots is not 0 i.e. q not equals 0) for a=1, b=0 => not possible so a= -1 and b =2 --> req condition i may be wrong somewhere though,,hmm you may also that p/q should be -1/2